Find the standard equation of the circle passing through a given point with a given center.
center (1,4)and passing through (-3,-2)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Find the standard equation of the circle passing through a given point with a given center.
center (1,4)and passing through (-3,-2)
-----------------
Find the radius squared
r^2 = x^2 + y^2
r^2 = 16+36 = 52
(x-1)^2 + (y-4)^2 = 52


Answer by stanbon(75887)

Find the standard equation of the circle passing through a given point with a given center.
center (1,4) and passing through (-3,-2)
---------------------------------------------
radius = distance from center to the point:
r = sqrt[(4--2)^2 + (1--3)^2] = sqrt[36 + 16] = sqrt(52)
---------------------------------------------------------------
Equation:
(x-1)^2 + (y-4)^2 = 52
============================
Cheers,
Stan H.